OpenStudy (rane):

A box of breakfast cereal has "contains 500g of breakfast cereal" printed on it. Suppose that in fact the weight of breakfast cereal contained in these boxes is normally distributed with a mean of 512g and a st. deviation of 8g. Determine the probability that a randomly chosen box of cereal contains less then 500g

OpenStudy (kropot72):

This can be solved by finding the z-score for 500g and then using a standard normal distribution table. Do you need help with these?

OpenStudy (kropot72):

The z-score can be found from the formula: \[\large z=\frac{X-\mu}{\sigma}\]

OpenStudy (rane):

X = the given weight or?

OpenStudy (kropot72):

X is the given value of the random variable. In this case X = 500g.

OpenStudy (kropot72):

\[\large z=\frac{X-\mu}{\sigma}=\frac{500-512}{8}=you\ can\ calculate\] After you calculate the z-score, you need to use a standard normal distribution table to find the required probability.

OpenStudy (rane):

1.5

OpenStudy (perl):

you can use a calculator TI 83 or TI 84

OpenStudy (perl):

normalcdf ( -1E99, 500, 512, 8 ) =

OpenStudy (kropot72):

Not really. What is the value of the numerator: 500 - 512 = ?

OpenStudy (perl):

.0668

OpenStudy (rane):

500-512 = -12/8 = -1.5

OpenStudy (perl):